Archive for July 2016

Next Meeting: Sunday 31st July at Petersfield community Centre

No Apiary meeting this Sunday  ( 31st July) as we will be extracting honey taken from the Apiary.    We have at least 5 or 6 supers to extract, so lots of help would be appreciated.

We will be using the new electric 12 frame  club spinner.   and tanking off into buckets to save time, and to let the honey settle.

Starting at 14:00  and should be finished by 17:00

Hope to see you all there.    (ps  I  have the key to the door this time)

Richard, Liz and Mel.  Apiary custodians.

Apiary Records – Sunday 24 July 2016

Weather breezy, warm & cloudy.  Starting to rain.  Identified supers ready for extraction next week (porter bee escapes to be added on Thursday).

Hive#1 – Not checked, poss virgin queen

Hive#2 – Not checked, poss virgin queen

Hive#4 – Inspected by Brad.  Brood on 9 frames, question whether sufficient room. 4 supers, 2 ready for extraction.  Swap brood frame for foundation at next inspection.  Queen could do with re-marking. Add super next week.  Some sac brood.

Hive#5 – Inspected by Mel.  Brood + half.  Brood on 1 frame in deep.  Stores and lots of pollen.  Supers not ready for extraction. Super added last week not started to be filled.  Feisty bees.

Hive#6 – not checked.

Hive#7  – inspected by Liz.   Searched for new queen but not seen.  Temper good so poss queenright.  Believe queen in bottom but QE under top deep.  Bottom super ready to extract next week.

Hive #7A – not checked.

Hive #8 – inspected by Richard.  Queen not seen. Brood on 7 frames.  Super not ready for extraction.  Plenty of room.

Hive#15F – inspected by Peter.  QC knocked down.  Queen seen.  Angry bees.  Second super added. Brood on 7 frames.

Hive#15H – inspected by Richard & Mel.  2 X deep (no supers last inspection).  Added super.  Top super ready to extract next week.  Moved a further 3 feet towards Hive 3 site.

Hive 16 B – inspected by Brad.  Very strong Nuc. Brood on 5 frames – no room.  Needs hiving.

Hive 16C – inspected by Richard.  Nuc.  Brood on 3 framnes. Needs super frames replacing with deeps after brood hatched.

 

Apiary Records: Tuesday 19th July

Was meant to be SBI  recheck on colonies to give clearance from EFB.   Due to SBI being off sick it was cancelled at the last minute,  but as we had delayed going through the colonies on Sunday we had to check them over.

Weather :-  HOT  and Sunny,  Little wind,  24 degrees.

Hive#1    not gone through as poss Virgin Queen

Hive#2    not gone through as poss virgin Queen

Hive#4    Inspected by Richard and Liz.   Brood Chamber:   National Deep.  QNS, no Qc’s. Brood and new eggs, room/stores ok.   4 supers,  3 ready to take off.  –  lots of play cups but none charged.  Temper 7

Hive#5   Inspected By Richard and Liz.    Brood chamber: Deep and Half.   QNS,   3 hatched Qc’s.  1 deep frame of brood/eggs . lot of drones – first time gone through for a month.  Temper 7

Hive#6  Not gone through as poss virgin Queen.

Hive#7  Not gone through

Hive# 7A    Inspected by  Richard and Liz,   Brood chamber:  4 Shallows ( Peter Street’s Method)                QNS,  Several play cups, not charged,  5 f bas in top super, room and stores ok. temper 8

Hive#8     Inspected by Richard and Liz.    Brood chamber: Deep.   QNS,  no Qc’s.  2 f BAS,  stores ok but room very limited,  Q laying right up to the frames .   2 supers, both near full, needs another next visit.  Temper  9.

Hive# 15 F,   Inspected by Richard and Liz,  Brood chamber: Deep.    QS,   1 QC,  Knocked down. 6 f BAS,  stores /room ok.  1 super( nearly full) –  Needs another Super next visit. Temper 9

Hive#15 H,    Inspected by Richard and Liz,  Brood chamber: Deep.   QNS,  3 Charged Qc’s knocked down. 6 F BAS with new eggs.  lack of room to lay so made up another Deep and moved another 3 ft towards Hive 3 slot.   Temper: 9.

Nuc’s  16B and 16C  not gone through.

Next Apiary Meeting: Sunday 24th July : 14:00 hrs, Still on EFB restrictions

Due to Debbie Burton, SBI being off sick we have still to wait for clearance for the apiary,  but we have taken every precaution suggested by the SBI and have culled the other two weak colonies that may have been likely to catch it, namely the Top Bar and Warre hives.                               The remaining colonies are progressing well,but due to the restrictions we have been unable to take any full supers off and they are now stacking up, having left them on the colonies.  No 4  now has 4 supers on, 3 are full and ready to be extracted.  We are working with the Regional bee inspector to try and get permission to take the full supers off.

Whatever the outcome, we need to go through the colonies to check on the usual 5 points.            1 –  any Queen cells.  2 – If Queen right,  eggs, brood etc.   3 – Room to expand, and for the queen to lay.   4 – Stores – enough food for the coming weeks.    5 – Disease,  any signs of DWV,  or EFB/AFB.

Lots to do so hopefully see as many as you can there to help out.    You can learn a lot from going through hives, other than your own and chat about how your bees are doing…                                   See you there.

Richard, Liz and Mel.     Apiary Custodians.

 

Apiary Record : 10th July 2016.

Still under EFB shutdown until being rechecked by the SBI on the 19th July.  So quick check through, making sure no colonies were about to swarm and that they had enough room,stores and that they had no sign of any disease.

Weather:   Sunny after earlier rain/drizzle.  19 degrees C,  Breezy.

Hive#1   Inspected by Liz,    Brood Chamber  14 x 12,  QNS,  No Qc’s,  No Brood, No eggs, stores/room ok.   1 super  75% full,   Queenless..  Introduced a frame with 2 QC’s from 15H( soon to be the new Colony 3).  Drone Trap removed,  far too many drones in the hive. leave 4 weeks.  Temper : 9

Hive#2   Not Checked as sealed QC left on f2 on 3/7/16    will be checked on 19/7 by SBI

Hive#4  Inspected by Liz.   Brood chamber -National Deep.  QNS,  2 Qc’s Brood  AS,  and new day old eggs,  2 Supers  nearly full, stores/room ok.   –  Knocked down both charged Qc’s.  Temper: 7

Hive#5  Not gone through as poss virgin Queen.

Hive#6  Inspected by Liz,    Brood chamber,  Deep and a Shallow.   QNS, possibly swarmed.        4 charged Qc ‘s   Brood  5D, 4S,  2 supers, top full, –  knocked down 2 Qc’s leaving 1 sealed, and 1 charged but open, on nframe with red pin.  No eggs seen at all.-  leave for 4 weeks.                   Temper:  6.5

Hive#7  Inspected by Liz,   Brood chamber: Deep.   QNS,  no Qc’s, stores/room ok,  2 supers.     new eggs in drone comb only, so may be a drone laying queen, Combined with Nuc 16A as 2 Sealed Qc’s on f5.

Hive#7A   Inspected by Richard,    Brood chamber:  4 Shallows.   QNS, no Qc’s.  7 frames of brood, all in top super. some 6 -7 day old larva,  no new eggs.  Q gone?  Dirty comb in bottom supers needs replacing.

Hive#8    Inspected by Richard,    Brood chamber : Deep.   QNS, 2 charged Qc’s,  6 f BAS, new eggs, 2 supers, top full.  stores/room ok,   Charged Qc’s taken down.  lots of bees looks like they may swarm.

Hive#15F  Inspected by Liz.     Brood chamber: Deep.  QS,  no Qc’s.  6 f BAS,  stores/room ok,     1 super,  –  Knocked down several play cups.

Hive#15H   Inspected by Richard.   Brood chamber: Deep.   QS,  2 QC’s ,  7 f BAS and new eggs.  Stores/room ok,  1 super . Part filled,-   Frame with 2 Qc’s given to Hive 1.    Temper:  9

Nuc # 16A   Combined with Hive#7.   but no Q found, 2 sealed QC’s  left.

Nuc# 16B   Not gone through – though bees bringing in pollen

Nuc# 16C  Not gone through –  bees flying

 

 

Next Apiary Meeting: Tuesday 19th July, 09:30 Debbie Burton SBI revisit.

No Apiary meeting on Sunday due to the SBI visit on Tuesday 19th,  All welcome to attend, but please note that it is in the morning at 09:30.   Club suites  and gloves only until we get the all clear.

Richard, Liz and Mel.    Apiary Custodians.

Apiary Record: 3 July 2016

Apiary continues to be in EFB lockdown.  6-week inspection due 10 July.

Hive #1 – not checked

Hive #2 – continues to set down charged queen cells.  9 frames of sealed brood, some older larvae but no eggs seen.  Removed 10 charged QCs, left 2.   Drawing top super, some capped stores.

Hive #4 – 9 frames of brood.  15+ play cups, 3 with eggs but not charged.  Big colony.  Some bald brood and sac brood.

Hive #5 – not checked.

Hive #6 – Took down 2 queen cells leaving 1 sealed 1 charged.  Marked frame with red pin.  No eggs.   Brood on 5 deep frames and 4 shallows.

Hive #7 – 1 day old eggs in drone comb only.

Hive #7A –  Sealed brood but no eggs/larvae in any shallows below QX.      Large number of dead drones on top of QX.  Eggs/larvae/sealed brood in super.  Removed QX.  Bees grumpy.

Hive #8 – new eggs 1-3 days.  Calmed down and temper fine. Super half full.

Hive #15H – continues its weekly migration towards Hive #3 site.  Queen seen.  5 queen cells – all removed.  Queen seen.  Better temper.

Hive #15F – 6 frames of brood, 3 stores.  Whole frame of new eggs.  Good temper.

Hive #16A – Nuc created from Hive #4.  Queen found and marked white (NB one worker got in the way and also marked white!).  Transferred all frames to deep, plus frame of stores.  Some sealed cells opened (random pattern) and varroa seen.

Hive #16B – Nuc created from Hive #4.  Brood in all stages, healthy development.  Queen seen but disappeared before marker pen located.